OpenStudy (anonymous):

Solve for x. 4^(9x) = 9^(x+4) Write the exact answer using base-10 logarithms.

OpenStudy (amistre64):

9x log(4) = (x+4) log(9) log(4) x+4 ----- = ---- log(9) 9x

OpenStudy (amistre64):

the right goes to; (1/9) + (4/9x)

OpenStudy (amistre64):

log(4) 1 4 ----- - --- = --- log(9) 9 9x 4 9x = ----------------- (log(4)/log(9)) - (1\9) 4 x = ---------------------- 9 [(log(4)/log(9)) - (1\9)]

OpenStudy (amistre64):

maybe :)

OpenStudy (anonymous):

I see where you're going with this. But when I punched the thing into wolframalpha. I get a different answer. its x = 4 log(3) / (9log(2) - log(3))

OpenStudy (anonymous):

Do you have any idea how it came to that solution?

OpenStudy (amistre64):

wolfram prolly simplified it; or i just plain missed it :)

OpenStudy (amistre64):

log(3)?

OpenStudy (anonymous):

yes log(3)

OpenStudy (amistre64):

log(4) and log(9) are the only ones I see ..... cant tell why they got that

OpenStudy (anonymous):

yea that is why im confused :/

OpenStudy (amistre64):

thats the root; for x ... which might be the same thing have to look at

OpenStudy (anonymous):

ah okay thanks for the help :)

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=y+%3D+%284^%289x%29%29+and+y%3D9^%28x%2B4%29 is what it gives me

OpenStudy (amistre64):

same thing in exact form lol

OpenStudy (anonymous):

i think you inputted it wrong this is what i got http://www.wolframalpha.com/input/?i=4^%289x%29+%3D+9^%28x%2B4%29

OpenStudy (amistre64):

we got the same results ..... the log(3) stuff

OpenStudy (anonymous):

oh i didnt notice. lol

OpenStudy (anonymous):

so is that right or is the answer you gave me right? or are they the same?

OpenStudy (amistre64):

i cant see where mine is off; but i dint try to simplify nuthin; try thers, they seem to be smarter than me :)

OpenStudy (anonymous):

but I dont know how they got there >.<

OpenStudy (amistre64):

so they are the same :)